Among various ethnic groups, the standard deviation of heights is kno n to be approximately three inches we sh to construct 95% con dence interval for the mean he i of males from a certain country. Forty-eight males are surveyed from a particular country. The sample mean is 71 inches. The sample standard deviation is 2.3 inches D Part (a) Part (b) D Part (e) E Part (d) Construct a 95% confidence interval for the population mean height of males of this country. 0) State the oonfidence interval. (Round your answers to two decimal places) () Sketch the graph C.L. = ()Calculate the ermor bound. (Round your answer to two decimal places)
If you hand write your work please make it legible and clear as to which answer goes with what problem. If you know how do these with the calculator please write the steps. This question is with this chapter: Ch. 8 Confidence Intervals in Illowsky and Dean, "Introductory Statistics", ed.
Answer #1

Here confidence interval to get = 95%

Critical test statistic for sample size n = 48 is

as degree of freedom = 48 -1 = 47

t47 = 2.01174

sample mean \bar x = 71 inches

sample standard deviation s= 2.3 inches

Standard error of sample mean se0 = s/ sqrt(n) = 2.3/sqrt(48) = 0.3320

Margin of error = 0.3320 * 2.01174 = 0.66785

95% confidence interval = \bar x +- MOE = 71 +- 0.66785 = (70.33 inches, 71.67 inches)

Error bound = 0.67 inches

Here C.L. = 0.95

alpha/2 = 0.025

Here when we neeed calculator we need only one value which is critical value of t then go to "stat" => "Tests" -> T test and put df in that and two tailed test. then you will get the critical value.